# Bridging the Long-Term Gap: A Memory-Active Policy for Multi-Session Task-Oriented Dialogue

### v1

@@ B.3 Human Validation Protocol
B.4 Intent-driven QA Memory
For each historical session, we generated an intent description and the corresponding QA memory (as shown in Figure 13) for the objectives of that intent description. The QA memory consists of multiple QA pairs, where each query is a question about a task attribute under that intent, and the answer is the slot value corresponding to that task attribute.
B.5 Dataset Design Rationale
@@
Appendix C Supplementary Experimental Results
C.1 Memory Activation Comparision
Table 12 compares the performance of different activation modules on memory retrieval. text-embed3-small achieves the highest recall across all thresholds, with 0.702 at Recall@3, 0.792 at Recall@5, and 0.905 at Recall@10, demonstrating superior retrieval capability. Among other models, nv-embed-v2 and bge-large-en-v1.5 also perform well, while traditional retrieval methods like BM25 remain competitive at Recall@10 but lag behind embedding-based methods at lower recall levels. T5-base and BERT-based models exhibit lower recall, suggesting that general pre-trained models are less effective for specialized memory retrieval. These results highlight text-embed3-small as the most effective choice for long-term memory activation in multi-session dialogues.
C.2 Effectiveness of the Proactive Response Strategy
@@
C.3 Human Evaluation Details
Table 15 presents the results of human evaluation, including accuracy, informativeness, and coherency scores. Accuracy is rated on a scale of 0 to 1, while informativeness and coherency are rated from 0 to 3. The average scores in 5 are computed using a weighted sum with weights of 1, 1/3, and 1/3. All evaluations were conducted in a blind review manner to compare the response quality of w/o MAP and w/ MAP. Additionally, the Confirmation-type Response type assesses the response quality after memory-guided dialogue planning, while the multi-turn evaluation focuses on dialogues under the proactive response strategy, continuing until task completion or forced termination.
@@ C.3 Human Evaluation Details
LLaMA3-8B w/o MAP 0.62 10.47 28.59
w/ MAP 0.56 9.86 30.39
[…]
w/ MAP 0.55 10.90 31.28
[…]
w/ MAP 0.56 6.66 24.64
[…]
w/ MAP 0.68 13.6 35.20
@@
GPT4o-mini w/o MAP 0.62 1.83 1.90 0.81 1.92 2.44
w/ MAP 0.65 2.38 2.48 0.87 1.93 2.74
[…]
w/ MAP 0.61 1.98 2.16 0.88 2.51 2.71
[…]
w/ MAP 0.54 1.70 2.30 0.92 1.93 2.47
[…]
w/ MAP 0.61 2.06 2.08 0.93 2.74 2.85
@@
Table 15: Comparison of different models on human evaluation metrics: accuracy, informativeness, and coherence. The results are presented for both confirmation-type responses and multi-turn dialogue settings, comparing standard inference (‘w/o MAP‘) with memory-augmented processing (‘w/ MAP‘).
C.4 Additional Evaluation Metrics
@@
Table 13 compares the performance of task-oriented dialogue models with and without memory-augmented processing (MAP) across Slot Accuracy, BLEU, and ROUGE metrics. The results reveal a trade-off between structured slot accuracy and response fluency. In most models, MAP slightly reduces slot accuracy, as seen in LLaMA3-8B, which drops from 0.62 to 0.56, and Mistral-7B, which decreases from 0.59 to 0.56. However, GPT4o-mini benefits from MAP, achieving the highest slot accuracy of 0.68. BLEU scores generally decline, suggesting that MAP shifts responses away from verbatim accuracy towards greater contextual adaptability. Mistral-7B drops from 10.90 to 6.66, and LLaMA3-8B decreases from 10.47 to 9.86. Conversely, ROUGE scores improve with MAP in several cases. LLaMA3-8B increases from 28.59 to 30.39, and Qwen-7B rises from 29.77 to 31.28, indicating enhanced informativeness and coherence. However, Mistral-7B experiences a slight decrease in ROUGE from 28.42 to 24.64. Overall, the results suggest that MAP enhances response informativeness while slightly compromising slot accuracy and BLEU, highlighting a trade-off between structured information retention and more natural, contextually aware responses.
@@
Table 14 presents the performance comparison between AutoTOD and MAP on Slot Accuracy, BLEU, and ROUGE. The results indicate that MAP consistently outperforms AutoTOD across all three metrics, demonstrating its effectiveness in enhancing dialogue quality. Slot Accuracy improves from 0.61 to 0.68, indicating better tracking of task-specific information. BLEU increases from 3.34 to 5.47, reflecting more precise and fluent responses. ROUGE also shows a slight improvement, rising from 24.07 to 25.03, suggesting that MAP enhances informativeness and coherence. These results highlight the advantages of memory-augmented processing, which enables more accurate and contextually relevant dialogue generation.
Appendix D Case Study Detail
D.1 Multi-session Dialogue Context Comparison
@@
Figure 14 presents four different configurations of conversation contexts not shown in the main paper. Specifically, (1) Full conversation history includes every session from the dialogue history as prompt input to the reader. (2) Retrieval-based methods retrieve the dialogue sessions most relevant to the current session (Session 23) and append them to the reader’s context (3) Retrieving a summary compiles a summary of past sessions (Sessions 1 to 22) for inclusion alongside the current context. Finally, (4) MAP integrates QA memory with the Session 23 context to generate responses. By illustrating these detailed contexts, Figure 14 provides further insights into how each approach manages multi-session dialogue.
D.2 MAP vs. RAG
@@
To better understand how CoT reasoning and memory reranking affect confirmation response generation, we present a step-by-step case study comparing MAP and standard RAG (Appendix Table 16). In this example, the user attempts to confirm a restaurant reservation. While both systems retrieve similar QA memory candidates, the standard RAG model fails to detect missing slot information (e.g., number of people), resulting in an incomplete and partially inaccurate response. In contrast, MAP use Chain-of-Thought explicitly identifies missing task information (e.g., time, headcount) through reasoning, refines the retrieved memories via the Memory Judger, and generates a more complete and contextually appropriate confirmation. This illustrates how structured reasoning and selective memory grounding improve slot coverage and reduce factual errors in multi-turn dialogue.
Figure 6: Impact of memory judger on memory activation performance across different embedding models.
@@
Prompts of the Dataset Generation User Prompt: """ Help me generate an English conversation under the {dialogue_intent} intent, where {task_goal}. The conversation should be between a user and an assistant, and it should be split into {task_goal_length} sessions at different points in time, with continuity and connection between the sessions and each session should not less than 6 turns. Additionally, the final session must include a assistant response containing a complete confirmation-type utterance before the user confirms, and this utterance should be marked with ‘is_confirmation‘ set to ‘True‘. and the user must provide a final confirmation response at the end of the final session. For all other sessions, the conversation should end with an assistant’s polite declarative statement. """ System Prompt: """ You are dialogue generator assistant. The sessions should be clearly separated, and the conversation should be formatted as follows: Each turn should be a dictionary entry. The conversation should be in the format of a list of sessions, where each session is a list of dictionaries representing each turn. Each dictionary entry should have two keys: speaker (either ’user’ or ’assistant’) and text (the spoken dialogue). Except for final session, each session should be a seperate dialogue and include a complete dialogue structure, beginning with a greeting from the user and ending with an assistant’s polite declarative statement. Feel free to expand the dialogue with additional relevant details, but avoid redundant expressions or repeating the same phrases. Reponse me with a json format { "sessions": [ [ { "speaker": "xx", "text": "xx" }, { "speaker": "xx", "text": "xx" } ] ] } """
Figure 7: Prompts of the Dataset Generation
@@
Prompts of the Task Slot Querying Generation """ Please help me generate questions, based on the provided {conversation history}, that correspond to unanswered attributes in the task goal {task_attributes}. 1. The questions should start with ’What,’ ’When,’ ’Why,’ ’How,’ or ’Where.’ 2. Ensure that the generated questions are in third person. fill the following json: { [Question], } """
Figure 8: Prompts of the Task Slot Querying Generation
@@
Prompts of Confirmation Response Generation """ You are an dialogue assistant. Generate a confirmation response based on the userś utterance. Include any relevant task goals [TASK GOALS] identified in the dialogue or related memory [MEMORY]. If [MEMORY] is unavailable, construct your response accurately and comprehensively using the provided conversation details. Ensure your reply acknowledges the userś request clearly and incorporates relevant information from both the dialogue and the related memory units [MEMORY]. [[[[TASK GOAL]]]] {task_goal} [MEMORY] {memory} """
Figure 9: Prompt of Confirmation Response Generation
@@
Prompt of Dialogue State Tracking on MultiWOZ 2.2 """Consider the following list of concepts, called "slots" provided to you as a json list. "slots": { "attraction-area", "attraction-name", "attraction-type", "bus-day", "bus-departure", "bus-destination", "bus-leaveat", "hospital-department", "hotel-area", "hotel-bookday", "hotel-bookpeople", "hotel-bookstay", "hotel-internet", "hotel-name", "hotel-parking", "hotel-pricerange", "hotel-stars", "hotel-type", "restaurant-area", "restaurant-bookday", "restaurant-bookpeople", "restaurant-booktime", "restaurant-food", "restaurant-name", "restaurant-pricerange", "taxi-arriveby", "taxi-departure", "taxi-destination", "taxi-leaveat", "train-arriveby", "train-bookpeople", "train-day", "train-departure", "train-destination", "train-leaveat", } Now consider the following dialogue between two parties called the "system" and "user". Can you tell me which of the "slots" were updated by the "user" in its latest response to the "system"? Present the updates in JSON format. If no "slots" were updated, return an empty JSON list. If you encounter "slots" that were requested by the "user" then fill them with "?". If the user informed that he did not care about a "slot", fill it with "dontcare". Return the output in JSON format and no explanation! {dialogue} """
Figure 10: Prompt of Dialogue State Tracking on MultiWOZ 2.2
@@
Prompts of GPT4 Evaluation """ You are a strict and objective evaluator. Your task is to assess the quality of the final predicted response using the provided conversation context, the user’s target goal attributes, and a reference answer. Your evaluation should be fair, professional, and reflect an expert judgment of the response’s quality. [Dialogue Context] $\{\{conversation_{h}istory\}\}$ [Task Goal] $\{\{task\_goal\}\}$ [reference_answer] $\{\{reference\_anwser\}\}$ [predict_answer] $\{\{predict\_answer\}\}$ Evaluation Criteria: Requirement Alignment: Does the final predict_answer meet the user’s task goal? Content Accuracy: Is the information in the final response correct, clear, and logically organized? Language Quality: Is the language fluent, coherent, and readable? Are there any obvious grammatical or word choice errors? Comparison to Reference Answer: Compared to the reference answer, how does the final response differ in terms of completeness, professionalism, and clarity? Overall Score: Assign a score from 1 to 10 (10 being the best), considering all of the above factors. The evaluation must be structured in the following JSON format: “‘json { "Rationale": "<Explain the rationale of your score.>", "Score": <An integer score from 1 to 10.> } “‘ """
Figure 11: Prompts of GPT4 Evaluation
@@
MS-TOD dialogue session structure { "session_id": 9, "domain": "Travel", "reference_dialogue_id": "66_00101", "exist_confirmation": true, "intent": "ReserveHotel", "content": [ { "speaker": "user", "utterance": "Hi again, Im ready to finalize the booking for Aloft Portland Airport At Cascade Station.", "is_confirmation": false }, { "speaker": "assistant", "utterance": "Just to confirm, you are booking 1 room at Aloft Portland Airport At Cascade Station, Portland, from March 5th to March 7th. The room is a standard king room with free Wi-Fi and a 24-hour cancellation policy. Is that correct?", "is_confirmation": true }, { "speaker": "user", "utterance": "Yes, that is correct.", "is_confirmation": false }, { "speaker": "assistant", "utterance": "Excellent! Your room has been successfully booked. You will receive a confirmation email shortly.", "is_confirmation": false }, { "speaker": "user", "utterance": "Thank you so much for your help!", "is_confirmation": false }, { "speaker": "assistant", "utterance": "Youre welcome! Have a great stay in Portland.", "is_confirmation": false } ], "task_goal": { "hotel_name": "Aloft Portland Airport At Cascade Station", "location": "Portland", "check_in_date": "March 5th", "check_out_date": "March 7th", "number_of_rooms": 1 } }
Figure 12: MS-TOD Session Structure.
@@
MS-TOD Intent Description and QA Memory { "9": { "intent_description": "The users intent is to finalize and confirm a hotel booking for a specific room at Aloft Portland Airport At Cascade Station, including details about the stay dates and room type.", "qa_summary": [ { "Question": "What type of room did the user book?", "Answer": "The user booked a standard king room." }, { "Question": "When is the users reservation?", "Answer": "The users reservation is from March 5th to March 7th." }, { "Question": "Where is the users reservation located?", "Answer": "The users reservation is located at Aloft Portland Airport At Cascade Station." }, { "Question": "What amenities are included in the users reservation?", "Answer": "The users reservation includes free Wi-Fi." }, { "Question": "What is the cancellation policy for the users booking?", "Answer": "The cancellation policy for the users booking is 24 hours." } ] } }
Figure 13: Intent description and QA Memory in MT-TOD.
Process MAP RAG
@@
Memory Judger (Reranking)
@@
Table 16: Step-by-step comparison of MAP vs. standard RAG in confirmation response generation.
Figure 14: Comparison of confirmation response generation across four approaches: (1) Direct Prompting with the full conversation history, (2) Hybrid RAG retrieving relevant dialogue history, (3) Hybrid RAG retrieving a summary of the conversation, and (4) MAP with intention-based QA memory.

### v2

@@ Appendix B Appendix B
Intent-driven QA Memory
For each historical session, we generated an intent description and the corresponding QA memory (as shown in Figure A.7) for the objectives of that intent description. The QA memory consists of multiple QA pairs, where each query is a question about a task attribute under that intent, and the answer is the slot value corresponding to that task attribute.
Dataset Design Rationale
@@
Appendix C Appendix C
Memory Retrieval Comparision
Table 11 compares the performance of different retrieval methods. text-embed3-small achieves the highest recall across all thresholds, with 0.702 at Recall@3, 0.792 at Recall@5, and 0.905 at Recall@10, demonstrating superior retrieval capability. Among other models, nv-embed-v2 and bge-large-en-v1.5 also perform well, while traditional retrieval methods like BM25 remain competitive at Recall@10 but lag behind embedding-based methods at lower recall levels. T5-base and BERT-based models exhibit lower recall, suggesting that general pre-trained models are less effective for specialized memory retrieval. These results highlight text-embed3-small as the most effective choice for long-term memory activation in multi-session dialogues.
Effectiveness of the Proactive Response Strategy
@@ Appendix C Appendix C
Human Evaluation Details
Table 14 presents the results of human evaluation, including accuracy, informativeness, and coherency scores. Accuracy is rated on a scale of 0 to 1, while informativeness and coherency are rated from 0 to 3. The average scores in Table 3 are computed using a weighted sum with weights of 1, 1/3, and 1/3. All evaluations were conducted in a blind review manner to compare the response quality of FCP and MemGuide. Additionally, the Confirmation-type Response type assesses the response quality after memory-guided dialogue planning, while the multi-turn evaluation focuses on dialogues under the proactive response strategy, continuing until task completion or forced termination.
@@
LLaMA3-8B FCP 0.62 10.47 28.59
MemGuide 0.56 9.86 30.39
Qwen-7B FCP 0.48 10.33 29.77
MemGuide 0.55 10.90 31.28
Mistral-7B FCP 0.59 10.09 28.42
MemGuide 0.56 6.66 24.64
GPT4o-mini FCP 0.61 20.30 43.49
MemGuide 0.68 13.6 35.20
@@
GPT4o-mini FCP 0.62 1.83 1.90 0.81 1.92 2.44
MemGuide 0.65 2.38 2.48 0.87 1.93 2.74
LLaMA FCP 0.56 1.47 1.74 0.78 1.64 2.36
MemGuide 0.61 1.98 2.16 0.88 2.51 2.71
Qwen FCP 0.43 1.24 1.85 0.82 1.60 2.02
MemGuide 0.54 1.70 2.30 0.92 1.93 2.47
Mistral FCP 0.58 1.63 1.99 0.89 2.49 2.72
MemGuide 0.61 2.06 2.08 0.93 2.74 2.85
@@
Table 14: Comparison of different models on human evaluation metrics: accuracy, informativeness, and coherence. The results are presented for both confirmation-type responses and multi-turn dialogue settings, comparing standard inference (‘FCP‘) with memory-augmented processing (‘MemGuide‘).
Additional Evaluation Metrics
@@
Table 12 compares the performance of task-oriented dialogue models with and without memory-augmented processing (MemGuide) across Slot Accuracy, BLEU, and ROUGE metrics. The results reveal a trade-off between structured slot accuracy and response fluency. In most models, MemGuide slightly reduces slot accuracy, as seen in LLaMA3-8B, which drops from 0.62 to 0.56, and Mistral-7B, which decreases from 0.59 to 0.56. However, GPT4o-mini benefits from MemGuide, achieving the highest slot accuracy of 0.68. BLEU scores generally decline, suggesting that MemGuide shifts responses away from verbatim accuracy towards greater contextual adaptability. Mistral-7B drops from 10.90 to 6.66, and LLaMA3-8B decreases from 10.47 to 9.86. Conversely, ROUGE scores improve with MemGuide in several cases. LLaMA3-8B increases from 28.59 to 30.39, and Qwen-7B rises from 29.77 to 31.28, indicating enhanced informativeness and coherence. However, Mistral-7B experiences a slight decrease in ROUGE from 28.42 to 24.64. Overall, the results suggest that MemGuide enhances response informativeness while slightly compromising slot accuracy and BLEU, highlighting a trade-off between structured information retention and more natural, contextually aware responses.
@@
Table 13 presents the performance comparison between AutoTOD and MemGuide on Slot Accuracy, BLEU, and ROUGE. The results indicate that MemGuide consistently outperforms AutoTOD across all three metrics, demonstrating its effectiveness in enhancing dialogue quality. Slot Accuracy improves from 0.61 to 0.68, indicating better tracking of task-specific information. BLEU increases from 3.34 to 5.47, reflecting more precise and fluent responses. ROUGE also shows a slight improvement, rising from 24.07 to 25.03, suggesting that MemGuide enhances informativeness and coherence. These results highlight the advantages of memory-augmented processing, which enables more accurate and contextually relevant dialogue generation.
Appendix D Appendix D
Multi-session Dialogue Context Comparison
@@ Appendix D Appendix D
Figure LABEL:f-case_study presents four different configurations of conversation contexts not shown in the main paper. Specifically, (1) Full conversation history includes every session from the dialogue history as prompt input to the reader. (2) Retrieval-based methods retrieve the dialogue sessions most relevant to the current session (Session 23) and append them to the reader’s context (3) Retrieving a summary compiles a summary of past sessions (Sessions 1 to 22) for inclusion alongside the current context. Finally, (4) MemGuide integrates QA memory with the Session 23 context to generate responses. By illustrating these detailed contexts, Figure LABEL:f-case_study provides further insights into how each approach manages multi-session dialogue.
MemGuide vs. RAG
@@
To better understand how CoT reasoning and memory reranking affect confirmation response generation, we present a step-by-step case study comparing MemGuide and standard RAG (Appendix Table 15). In this example, the user attempts to confirm a restaurant reservation. While both systems retrieve similar QA memory candidates, the standard RAG model fails to detect missing slot information (e.g., number of people), resulting in an incomplete and partially inaccurate response. In contrast, MemGuide use Chain-of-Thought explicitly identifies missing task information (e.g., time, headcount) through reasoning, refines the retrieved memories via the Memory Judger, and generates a more complete and contextually appropriate confirmation. This illustrates how structured reasoning and selective memory grounding improve slot coverage and reduce factual errors in multi-turn dialogue.
@@
Prompts of the Dataset Generation User Prompt: ””” Help me generate an English conversation under the {dialogue_intent} intent, where {task_goal}. The conversation should be between a user and an assistant, and it should be split into {task_goal_length} sessions at different points in time, with continuity and connection between the sessions and each session should not less than 6 turns. Additionally, the final session must include a assistant response containing a complete confirmation-type utterance before the user confirms, and this utterance should be marked with ‘is_confirmation‘ set to ‘True‘. and the user must provide a final confirmation response at the end of the final session. For all other sessions, the conversation should end with an assistant’s polite declarative statement. ””” System Prompt: ””” You are dialogue generator assistant. The sessions should be clearly separated, and the conversation should be formatted as follows: Each turn should be a dictionary entry. The conversation should be in the format of a list of sessions, where each session is a list of dictionaries representing each turn. Each dictionary entry should have two keys: speaker (either ’user’ or ’assistant’) and text (the spoken dialogue). Except for final session, each session should be a seperate dialogue and include a complete dialogue structure, beginning with a greeting from the user and ending with an assistant’s polite declarative statement. Feel free to expand the dialogue with additional relevant details, but avoid redundant expressions or repeating the same phrases. Reponse me with a json format 1{ 2 "sessions": [ 3 [ 4 { 5 "speaker": "xx", 6 "text": "xx" 7 }, 8 { 9 "speaker": "xx", 10 "text": "xx" 11 } 12 ] 13 ] 14} ”””
Figure A.1: Prompts of the Dataset Generation
@@
Prompts of the Task Slot Querying Generation ””” Please help me generate questions, based on the provided {conversation history}, that correspond to unanswered attributes in the task goal {task_attributes}. 1. The questions should start with ’What,’ ’When,’ ’Why,’ ’How,’ or ’Where.’ 2. Ensure that the generated questions are in third person. fill the following json: { [Question], } ”””
Figure A.2: Prompts of the Task Slot Querying Generation
@@
Prompts of Confirmation Response Generation ””” You are an dialogue assistant. Generate a confirmation response based on the userś utterance. Include any relevant task goals [TASK GOALS] identified in the dialogue or related memory [MEMORY]. If [MEMORY] is unavailable, construct your response accurately and comprehensively using the provided conversation details. Ensure your reply acknowledges the userś request clearly and incorporates relevant information from both the dialogue and the related memory units [MEMORY]. [[TASK GOAL]] {task_goal} [MEMORY] {memory} ”””
Figure A.3: Prompt of Confirmation Response Generation
@@
Prompt of Dialogue State Tracking on MultiWOZ 2.2 ”””Consider the following list of concepts, called ”slots” provided to you as a json list. ”slots”: { ”attraction-area”, ”attraction-name”, ”attraction-type”, ”bus-day”, ”bus-departure”, ”bus-destination”, ”bus-leaveat”, ”hospital-department”, ”hotel-area”, ”hotel-bookday”, ”hotel-bookpeople”, ”hotel-bookstay”, ”hotel-internet”, ”hotel-name”, ”hotel-parking”, ”hotel-pricerange”, ”hotel-stars”, ”hotel-type”, ”restaurant-area”, ”restaurant-bookday”, ”restaurant-bookpeople”, ”restaurant-booktime”, ”restaurant-food”, ”restaurant-name”, ”restaurant-pricerange”, ”taxi-arriveby”, ”taxi-departure”, ”taxi-destination”, ”taxi-leaveat”, ”train-arriveby”, ”train-bookpeople”, ”train-day”, ”train-departure”, ”train-destination”, ”train-leaveat”, } Now consider the following dialogue between two parties called the ”system” and ”user”. Can you tell me which of the ”slots” were updated by the ”user” in its latest response to the ”system”? Present the updates in JSON format. If no ”slots” were updated, return an empty JSON list. If you encounter ”slots” that were requested by the ”user” then fill them with ”?”. If the user informed that he did not care about a ”slot”, fill it with ”dontcare”. Return the output in JSON format and no explanation! {dialogue} ”””
Figure A.4: Prompt of Dialogue State Tracking on MultiWOZ 2.2
@@
Prompts of GPT4 Evaluation ””” You are a strict and objective evaluator. Your task is to assess the quality of the final predicted response using the provided conversation context, the user’s target goal attributes, and a reference answer. Your evaluation should be fair, professional, and reflect an expert judgment of the response’s quality. [Dialogue Context] $\{\{conversation_{h}istory\}\}$ [Task Goal] $\{\{task\_goal\}\}$ [reference_answer] $\{\{reference\_anwser\}\}$ [predict_answer] $\{\{predict\_answer\}\}$ Evaluation Criteria: Requirement Alignment: Does the final predict_answer meet the user’s task goal? Content Accuracy: Is the information in the final response correct, clear, and logically organized? Language Quality: Is the language fluent, coherent, and readable? Are there any obvious grammatical or word choice errors? Comparison to Reference Answer: Compared to the reference answer, how does the final response differ in terms of completeness, professionalism, and clarity? Overall Score: Assign a score from 1 to 10 (10 being the best), considering all of the above factors. The evaluation must be structured in the following JSON format: “‘json { ”Rationale”: ”¡Explain the rationale of your score.¿”, ”Score”: ¡An integer score from 1 to 10.¿ } “‘ ”””
Figure A.5: Prompts of GPT4 Evaluation
@@
MS-TOD dialogue session structure 1{ 2 "session_id": 9, 3 "domain": "Travel", 4 "reference_dialogue_id": "66_00101", 5 "exist_confirmation": true, 6 "intent": "ReserveHotel", 7 "content": [ 8 { 9 "speaker": "user", 10 "utterance": "Hi again, Im ready to finalize the booking for Aloft Portland Airport At Cascade Station.", 11 "is_confirmation": false 12 }, 13 { 14 "speaker": "assistant", 15 "utterance": "Just to confirm, you are booking 1 room at Aloft Portland Airport At Cascade Station, Portland, from March 5th to March 7th. The room is a standard king room with free Wi-Fi and a 24-hour cancellation policy. Is that correct?", 16 "is_confirmation": true 17 }, 18 { 19 "speaker": "user", 20 "utterance": "Yes, that is correct.", 21 "is_confirmation": false 22 }, 23 { 24 "speaker": "assistant", 25 "utterance": "Excellent! Your room has been successfully booked. You will receive a confirmation email shortly.", 26 "is_confirmation": false 27 }, 28 { 29 "speaker": "user", 30 "utterance": "Thank you so much for your help!", 31 "is_confirmation": false 32 }, 33 { 34 "speaker": "assistant", 35 "utterance": "Youre welcome! Have a great stay in Portland.", 36 "is_confirmation": false 37 } 38 ], 39 "task_goal": { 40 "hotel_name": "Aloft Portland Airport At Cascade Station", 41 "location": "Portland", 42 "check_in_date": "March 5th", 43 "check_out_date": "March 7th", 44 "number_of_rooms": 1 45 } 46}
Figure A.6: MS-TOD Session Structure.
@@
MS-TOD Intent Description and QA Memory 1{ 2 "9": { 3 "intent_description": "The users intent is to finalize and confirm a hotel booking for a specific room at Aloft Portland Airport At Cascade Station, including details about the stay dates and room type.", 4 "qa_summary": [ 5 { 6 "Question": "What type of room did the user book?", 7 "Answer": "The user booked a standard king room." 8 }, 9 { 10 "Question": "When is the users reservation?", 11 "Answer": "The users reservation is from March 5th to March 7th." 12 }, 13 { 14 "Question": "Where is the users reservation located?", 15 "Answer": "The users reservation is located at Aloft Portland Airport At Cascade Station." 16 }, 17 { 18 "Question": "What amenities are included in the users reservation?", 19 "Answer": "The users reservation includes free Wi-Fi." 20 }, 21 { 22 "Question": "What is the cancellation policy for the users booking?", 23 "Answer": "The cancellation policy for the users booking is 24 hours." 24 } 25 ] 26 } 27}
Figure A.7: Intent description and QA Memory in MT-TOD.
Process MemGuide RAG
@@
Memory Filter (Reranking)
@@
Table 15: Step-by-step comparison of MemGuide vs. standard RAG in confirmation response generation.